OpenStudy (anonymous):

create a perfect square trinomial in order to solve: x²+3x+1=0

OpenStudy (john_es):

You must complete the square. \[x^2+2\frac{3}{2}x+\left(\sqrt{\frac{3}{2}}\right)^2+1-\left(\sqrt{\frac{3}{2}}\right)^2=0\] \[\left(x+\sqrt{\frac{3}{2}}\right)^2-1/3=0\] And solve.

OpenStudy (anonymous):

ah okay thanks, got it!

OpenStudy (john_es):

Um, it is incorrect. The correct form should be, \[x^2+2\frac{3}{2}x+\frac{9}{4}+1-\frac{9}{4}=0\\ \left(x+\frac{3}{2}\right)^2-5/4=0\Rightarrow x+3/2=\pm\sqrt{5}/2\Rightarrow\] \[x=-\frac{3}{2}\pm\frac{\sqrt{5}}{2}\]